MAT 122, OVERVIEW OF CALCULUS, SUNY AT STONY BROOK, FALL 2002

Lecturer: Prof. Christopher Bishop, bishop@math.sunysb.edu, office is 4-112 in Math Building.

Text: Applied Calculus, 2nd edition, Hughes-Hallet, Gleason, et. al.. Syllabus: We will cover approximately Chapters 1 to 7 with omissions

of certain sections. A more detailed syllabus will be provided.

Classdays. Classes begin on Tuesday, September 3 and end on Friday, December 13.

Holidays: There are no classes on Yom Kipur (Monday, September 16) and Thanksgiving (November 27-30).

Lecture times and locations:

Lecture Recitation day  time  instructor  room
01  Tu-Th 9:50-11:10 Bishop Old Eng 145
01  Th 2:20-3:15 Namazi  Phy P117
02  M  3:20-4:15 Savelev  Phy P117
03  W 9:25-10:20 Savelev  Phy P117
04  Tu  2:20-3:15 Namazi  Phy P117
02  MW 5:30-6:50 Y. Kim  Phy P112

There is no recitation for Lecture 2. Contact your instructors:

Instructor Office Office hours  email
Bishop  4-112 Tu 9-9:40, Th 12:40-1:20 bishop@math.sunysb.edu

Kim  2-119 Mon 3-4  ydkim@math.sunysb.edu
Savelyev  yasha@math.sunysb.edu
Namazi  hossein@math.sunysb.edu

Homework. A list of problems from the text and supplementary material will be handed out. The recitation instructor will assign, collect and return homework. He may also give weekly quizzes based on the assigned problems.

Examinations. There will be two midterm tests: Monday, October 14 and Tuesday, November 12. These will be given at night (8:30pm-10:10pm) in a location which will be annouced in class. The final exam is on Tuesday, December 17 at 11:00-1:30. The location will be announced in class during the last week of the semester.

Grades. Grades are computed using: Recitation 10%, Midterms 25% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: (Physics A-125/127) This is located on the A level of the physics building and provides help with pre-calculus material. The same room also contains the Calculus Resource Room where you can go for help with Calculus classes. It is staffed most days and some evenings; an exact schedule will be posted outside the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathe-

maticians, try http://www-groups.dcs.st-and.ac.uk/~history and click

on History Topics or the Biography Index. You can find information about math classes at Stony Brook at www.math.sunysb.edu

Drop dates: Last day for Add/Drop is Tuesday, September 17. Last day to switch to lower numbered math course is Monday, October 20. The last day to withdraw from a course is Friday, November 8.